# Towards reliable respiratory disease diagnosis based on cough sounds and vision transformers

### v1

@@ Abstract
Recent advancements in deep learning techniques have sparked performance boosts in various real-world applications including disease diagnosis based on multi-modal medical data. Cough sound data-based respiratory disease (e.g., COVID-19 and Chronic Obstructive Pulmonary Disease) diagnosis has also attracted much attention. However, existing works usually utilise traditional machine learning or deep models of moderate scales. On the other hand, the developed approaches are trained and evaluated on small-scale data due to the difficulty of curating and annotating clinical data on scale. To address these issues in prior works, we create a unified framework to evaluate various deep models from lightweight Convolutional Neural Networks (e.g., ResNet18) to modern vision transformers and compare their performance in respiratory disease classification. Based on the observations from such an extensive empirical study, we propose a novel approach to cough-based disease classification based on both self-supervised and supervised learning on a large-scale cough data set. Experimental results demonstrate our proposed approach outperforms prior arts consistently on two benchmark datasets for COVID-19 diagnosis and a proprietary dataset for COPD/non-COPD classification with an AUROC of 92.5%.
@@ I Introduction
Respiratory diseases such as COVID-19 and COPD are widespread in populations and early diagnosis is imperative for disease control and human wellbeing. According to [1], COPD was the third fatal disease in the world in 2022 and is expected to remain the third cause of death through 2050 at the global level. The current diagnosis of COPD requires measurements of lung functions using clinical equipment in hospitals which is expensive and inconvenient for mass screening. As a result, a significant fraction of COPD patients are undiagnosed in China and are exposed to risks of developing into fatal cancers.
@@
Recently, Zhang et al. [11] released a series of respiratory acoustic foundation models which are based on the latest self-supervised learning and are pre-trained on large-scale unlabeled respiratory sound data. In addition, the pre-trained foundation models are evaluated on a curated list of downstream tasks for respiratory sound classification or regression. The benchmarking in this work compares several modern deep models originally designed for audio event classification and demonstrates the superiority of their proposed models which are pre-trained specifically on respiratory sound data. However, their released models are based on backbone models which are of moderate sizes compared with the most advanced ones for audio event classification or image classification.
@@
We conduct extensive empirical studies to answer these research questions and propose a novel approach to cough-based respiratory disease classification. The contributions of our work can be summarised as follows:
@@ II-A Audio event classification
According to the model architectures, existing approaches to audio event classification in the literature can be categorized into two groups: Convolutional Neural Networks (CNN) [13, 14, 15, 16] and transformers [17]. Whilst early works employing ImageNet-pretrained CNN models on audio spectrograms achieved competitive accuracy, they have been outperformed by transformer-based models. Audio Spectrogram Transformer (AST) [17] was one of the first to apply vision transformers to audio event classification and performed superior to their CNN counterparts. Koutini et al. [18] proposed PaSST series by evaluating different variants of tailored ViT models and efficient training strategies for audio event classification. These works vary from one another in their employed backbone architectures and audio data preprocessing but share a similar transfer learning framework in which the ImageNet-pre-trained models are fine-tuned on audio spectrogram data for downstream tasks.
@@ IV-A Dataset
Our private LucaCough dataset is challenging yet more clinically useful as the negative samples cover not only the coughs of healthy subjects but also the coughs of subjects diagnosed with various respiratory diseases which may share similar cough signatures to COPD cough sounds. Specifically, the negative samples in our dataset include those collected from patients diagnosed with asthma, upper/lower respiratory tract infection (URI/LRI), Bronchiectasis and other non-COPD respiratory diseases. The LucaCough dataset consists of cough data from 3000 subjects (272 COPD and 2728 non-COPD subjects) in the training subset and cough data from 651 subjects (66 COPD and 586 non-COPD) in the test set. The cough sounds are recorded via several different mobile phones to make the data more diverse and enable the generalisation capabilities of the developed models. The data are manually annotated by at least three clinical experts with the help of medical records and lung function test readings. Informed written consent was obtained from all participants. The study was approved by the Ruijin Hospital Shanghai Jiaotong University School of Medicine Ethics Committee (No. 2023-199) and registered at ClinicalTrials.gov (NCT06082791).
@@ IV-E Evaluation of Respiratory sound data pretrained models
We conduct experiments to evaluate models pre-trained on respiratory sound data. The models employed in these experiments include those released by Zhang et al. [11] and EAT models further trained on cough data by ourselves. Firstly, we apply the released pre-trained models to the LucaCough datasets and present the experimental results together with those from [11] on the other two datasets. Subsequently, we continue the pertaining of EAT models on a large-scale cough sound dataset. As a result, we have a variant of the pre-trained EAT-large model dubbed EAT-large-cough as it was further trained on cough data. The EAT-large-cough model is subsequently fine-tuned on downstream tasks to obtain the classification results and compare them with other respiratory sound data pre-trained models (i.e. the OPERA series). To make a fair comparison, we use the official code released by [11] to fine-tune the OPERA models on downstream tasks whilst only linear probing results were reported in the original paper [11].
The comparison results are shown in Table II from which several conclusions can be drawn. Firstly, EAT-large-cough performs the best on three tasks consistently. Particularly, it outperforms EAT-large on three tasks with the AUROC margins ranging from 0.9 to 3.3 percentage points. This proves that continuous pre-training of the model on cough data can further strengthen its capabilities of classifying diseases based on cough sound data. Secondly, the OPERA models, though pre-trained on respiratory sound data in a self-supervised learning manner, perform inferior not only to EAT-large-cough but also to the models pre-trained on general audio data (i.e. PaSST-S, EAT-base and EAT-large). This may be attributed to the fact of low model capacities or the lack of pre-training on large-scale audio data like AudioSet-2M.
@@ IV-F On the effectiveness of SAM
In this subsection, we focus on exploring the effectiveness of the SAM optimizer in our particular cough classification tasks. Specifically, we compare experimental results without and with the use of SAM optimizer during fine-tuning whilst keeping all other experimental settings the same. The comparative study is conducted on the LucaCough dataset with three representative pre-trained models: PaSST-S, EAT-base, EAT-large and EAT-large-cough. The empirical results are shown in Table III. The experimental results demonstrate the use of SAM improves the AUROC values consistently for all four models.
@@
EAT-large-cough (Ours) 91.7 92.5 $\uparrow$
@@ V Conclusions and Discussion
The use of pre-trained models, particularly those pre-trained on respiratory sound data, has shown significant benefits in enhancing the classification of cough sounds. The continuous pre-training of models on cough data, as evidenced by the performance of the EAT-large-cough model, further strengthens the model’s ability to discern disease signatures from cough sound data.

### v2

@@ Abstract
Recent advancements in deep learning techniques have sparked performance boosts in various real-world applications including disease diagnosis based on multi-modal medical data. Cough sound data-based respiratory disease (e.g., COVID-19 and Chronic Obstructive Pulmonary Disease) diagnosis has also attracted much attention. However, existing works usually utilise traditional machine learning or deep models of moderate scales. On the other hand, the developed approaches are trained and evaluated on small-scale data due to the difficulty of curating and annotating clinical data on scale. To address these issues in prior works, we create a unified framework to evaluate various deep models from lightweight Convolutional Neural Networks (e.g., ResNet18) to modern vision transformers and compare their performance in respiratory disease classification. Based on the observations from such an extensive empirical study, we propose a novel algorithm Cough Search for cough-based disease classification based on both self-supervised and supervised learning on a large-scale cough data set. Experimental results demonstrate our proposed approach outperforms prior arts consistently on two benchmark datasets for COVID-19 diagnosis and a proprietary dataset for COPD/non-COPD classification with an AUROC of 92.5%.
@@ I Introduction
Respiratory diseases such as COVID-19 and COPD are widespread in populations and early diagnosis is imperative for disease control and human well-being. According to [1], COPD was the third fatal disease in the world in 2022 and is expected to remain the third cause of death through 2050 at the global level. The current diagnosis of COPD requires measurements of lung functions using clinical equipment in hospitals which is expensive and inconvenient for mass screening. As a result, a significant fraction of COPD patients are undiagnosed in China and are exposed to risks of developing into fatal cancers.
@@
Recently, Zhang et al. [11] released a series of respiratory acoustic foundation models which are based on the latest self-supervised learning and are pre-trained on large-scale unlabeled respiratory sound data. In addition, the pre-trained foundation models are evaluated on a curated list of downstream tasks for respiratory sound classification or regression. The benchmarking in this work compares several modern deep models originally designed for audio event classification and demonstrates the superiority of their proposed models which are pre-trained specifically on respiratory sound data. However, their released models are based on backbone models of moderate sizes compared with the most advanced ones for audio event or image classification.
@@
We conduct extensive empirical studies to answer these research questions and propose a novel cough-based respiratory disease classification approach Cough Search. The contributions of our work can be summarised as follows:
@@ II-A Audio event classification
According to the model architectures, existing approaches to audio event classification in the literature can be categorized into two groups: Convolutional Neural Networks (CNN) [13, 14, 15, 16] and transformers [17]. Whilst early works employing ImageNet-pretrained CNN models on audio spectrograms achieved competitive accuracy, they have been outperformed by transformer-based models. Audio Spectrogram Transformer (AST) [17] was one of the first to apply vision transformers to audio event classification and performed superior to their CNN counterparts. Koutini et al. [18] proposed the PaSST series by evaluating different variants of tailored ViT models and efficient training strategies for audio event classification. These works vary from one another in their employed backbone architectures and audio data preprocessing but share a similar transfer learning framework in which the ImageNet-pre-trained models are fine-tuned on audio spectrogram data for downstream tasks.
@@ IV-A Dataset
Our private LucaCough dataset is challenging yet more clinically useful as the negative samples cover not only the coughs of healthy subjects but also the coughs of subjects diagnosed with various respiratory diseases which may share similar cough signatures to COPD cough sounds. Specifically, the negative samples in our dataset include those collected from patients diagnosed with asthma, upper/lower respiratory tract infection (URI/LRI), Bronchiectasis and other non-COPD respiratory diseases. The LucaCough dataset consists of cough data from 3000 subjects (272 COPD and 2728 non-COPD subjects) in the training subset and cough data from 651 subjects (66 COPD and 586 non-COPD) in the test set. The cough sounds are recorded via several different mobile phones to make the data more diverse and enable the generalisation capabilities of the developed models. The data are manually annotated by at least three clinical experts with the help of medical records and lung function test readings. Informed written consent was obtained from all participants. The study was approved by the Ruijin Hospital Shanghai Jiaotong University School of Medicine Ethics Committee (No. 2023-199), Ruijin Hospital Luwan Branch Ethics Committee (2023-HXK-V1), Shanghai Jing’an District Central Hospital Ethics Committee (No. 2023-33), Shanghai Zhabei Central Hospital Ethics Committee (ZBLL2024030401001) and registered at ClinicalTrials.gov (NCT06082791).
@@ IV-E Evaluation of Respiratory sound data pretrained models
We conduct experiments to evaluate models pre-trained on respiratory sound data. The models employed in these experiments include those released by Zhang et al. [11] and EAT models further trained on cough data by ourselves. Firstly, we apply the released pre-trained models to the LucaCough datasets and present the experimental results together with those from [11] on the other two datasets. Subsequently, we continue the pertaining of EAT models on a large-scale cough sound dataset. As a result, we have a variant of the pre-trained EAT-large model dubbed Cough Search as it was further trained on cough data. The Cough Search model is subsequently fine-tuned on downstream tasks to obtain the classification results and compare them with other respiratory sound data pre-trained models (i.e. the OPERA series). To make a fair comparison, we use the official code released by [11] to fine-tune the OPERA models on downstream tasks whilst only linear probing results were reported in the original paper [11].
The comparison results are shown in Table II from which several conclusions can be drawn. Firstly, Cough Search performs the best on three tasks consistently. Particularly, it outperforms EAT-large on three tasks with the AUROC margins ranging from 0.9 to 3.3 percentage points. This proves that continuous pre-training of the model on cough data can further strengthen its capabilities of classifying diseases based on cough sound data. Secondly, the OPERA models, though pre-trained on respiratory sound data in a self-supervised learning manner, perform inferior not only to Cough Search but also to the models pre-trained on general audio data (i.e. PaSST-S, EAT-base and EAT-large). This may be attributed to the fact of low model capacities or the lack of pre-training on large-scale audio data like AudioSet-2M.
@@ IV-F On the effectiveness of SAM
In this subsection, we focus on exploring the effectiveness of the SAM optimizer in our particular cough classification tasks. Specifically, we compare experimental results without and with the use of SAM optimizer during fine-tuning whilst keeping all other experimental settings the same. The comparative study is conducted on the LucaCough dataset with three representative pre-trained models: PaSST-S, EAT-base, EAT-large and Cough Search. The empirical results are shown in Table III. The experimental results demonstrate the use of SAM improves the AUROC values consistently for all four models.
@@
Cough Search (Ours) 91.7 92.5 $\uparrow$
@@ V Conclusions and Discussion
The use of pre-trained models, particularly those pre-trained on respiratory sound data, has shown significant benefits in enhancing the classification of cough sounds. The continuous pre-training of models on cough data, as evidenced by the performance of the Cough Search model, further strengthens the model’s ability to discern disease signatures from cough sound data.